OpenStudy (anonymous):

If sinx = 5/13 and x is in quadrant 1, then tan(x/2) =

OpenStudy (anonymous):

first of all:sinx=[2tan(x/2)]/(1+tan^2(x/2))=5/13 =>5+5tan^2(x/2)=26tan(x/2) =>5tan^2(x/2)-26tan(x/2)+5=0 second of all:We note tan(x/2)=t =>5t^2-26t+5=0 delta=576=>sqrt(delta)=24 =>t1=5;t2=1/5; I. t1=5=>x/2=arctan5=>x=2((pi)n+ctg(5)),n is from Z; II. t2=1/5=>x/2=arctan1/5=>x=2((pi)n+ctg(1/5)),n is from Z;

OpenStudy (danielbarriosr1):

Deeeeeeeeeeeeeeeaaam

OpenStudy (anonymous):

 so what's cos x ? \[\tan \left( \frac{ x }{ 2 } \right)=\pm \sqrt{\frac{1-\cos x}{1+\cos x}}\] also, since x terminates in the first quadrant, we can assume \(0\le x\le \large\frac{\pi}{2}\) this means \(0\le \large\frac{x}{2}\le \large\frac{\pi}{4}\) so the tangent of x/2 should be positive.